OpenStudy (anonymous):

William traveled by scooter at an average speed of 62 miles per hour. Then, he traveled by car at an average speed of 36 miles per hour. In total he traveled 118 miles for 2 hours. How many minutes did William travel by car? ( Round to the nearest minute)

OpenStudy (anonymous):

distance = rate * time (ds=distance on scooter) (dc=distance on car) -------------------- ds = rs * ts dc = rc * tc (rc is rate of car; rs is rate of scooter; ts is time on scooter; tc is time on car)

OpenStudy (anonymous):

we already know rs and rc

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ds = 62 * ts dc = 36 * tc dt = total distance = 118 =62 * ts + 36 * tc tt = total time = 2 = ts + tc

OpenStudy (anonymous):

these are all of the equations we need, do you know how to solve for ts and tc. We have 2 equations and 2 unknowns at this point

OpenStudy (anonymous):

Wait, let me try.

OpenStudy (anonymous):

oh, also remember to convert hours to minutes once you have solved for tc

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you show me the steps. Im kind of stuck

OpenStudy (anonymous):

118 = 62 * ts + 36 * tc 2 = ts + tc we will use substitution

OpenStudy (anonymous):

ts = 2-tc

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

substitute (2-tc) wherever there is a ts in the following equation 118 = 62 * ts + 36 * tc becomes 118 = 62 * (2-tc) + 36 * tc

OpenStudy (anonymous):

can you solve this equation?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

start by distributing the 62 into the (2 - tc)

OpenStudy (anonymous):

124-62*ts

OpenStudy (anonymous):

i mean 124 - 62*tc

OpenStudy (anonymous):

124-26tc

OpenStudy (anonymous):

now put the rest of the equation around what you got

OpenStudy (anonymous):

-6 - 26tc

OpenStudy (anonymous):

118 = 124 - 26*tc + 36*tc Next step is to combine like terms

OpenStudy (anonymous):

oops my equation is wrong

OpenStudy (anonymous):

Yea i Was about to say that

OpenStudy (anonymous):

118 = 124 - 62*tc + 36*tc

OpenStudy (anonymous):

i think you just left out an equal sign

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

-6= -26tc

OpenStudy (anonymous):

Solve for tc. just remember to convert your answer to tc, which is in hours, into units of minutes (which is what the problem asks for).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 0.230769 repeating...

OpenStudy (anonymous):

Can you help me out?

OpenStudy (anonymous):

that fine, this number is in hours. we are interested in calculating the number of minutes spent in the car. Therefore, we need to convert hours to minutes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so we do 60/5

OpenStudy (anonymous):

and get 14 minutes?

OpenStudy (anonymous):

\[\frac{number of hours}{1}*\frac{60 minutes}{1hour}=numberofminutes\]

OpenStudy (anonymous):

your answer seems right

OpenStudy (anonymous):

Yes! Thank You!

OpenStudy (anonymous):

great!